1). 75 is 60 % of _____________ .
2). 36% of 900 is _____________ .

3). 45 is 30% of ______________ .

Answer by ikleyn(52777)

.
1). 75 is 60 % of _____________ .
2). 36% of 900 is _____________ .
3). 45 is 30% of ______________ .
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


(1)  75 is 60% of x.  Find x.


     To find x, write an equation, saying that 60% of x is 75;  then solve this equation.

         0.6x = 75  

            x = 75/0.6 = 125.


     ANSWER.  75 is 60% of 125.

Solved.


      
(2)  36% 0f 900 is  0.36*900 = 324.

     ANSWER.  36% of 900 is 324.

Solved.


(3)  45 is 30% of x.  Find x.


     To find x, write an equation, saying that 30% of x is 45;  then solve this equation.

         0.3x = 45  

            x = 45/0.3 = 150.


     ANSWER.  45 is 30% of 150.

Solved.


Everything is solved.

Notice that problem (3) is of the same type as problem (1).
Therefore, I solved them word-in-word.
The logic and the calculations are the same.
Only the numbers are different.

------------------------

To learn better,  wider and deeper,  see the lessons
    - Percentage problems
    - Percentage word problems (Type 1 problems, Finding the Part)
    - Percentage word problems (Type 2 problems, Finding the Rate)
    - Percentage word problems (Type 3 problems, Finding the Base)
in this site.


They are written specially for you.

Read them and become an expert in this area.

Start from the very first lesson of the list.

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.